OpenStudy (anonymous):

how do i start this. (by parts) find the integral of (t+5)(sqrt(4+6t))dt

OpenStudy (anonymous):

Set u = 4+6t and rewrite the integral in terms of u to make it easier to visualize the integration. Then you should get an integrand f(u)*sqrt(u) = f(u)*u^(1/2) and you can probably integrate that.

OpenStudy (anonymous):

You'd have to replace dt with the equation for du. Dunno if it will be easier.

zepdrix (zepdrix):

\[\large \int\limits (t+5)\sqrt{4+6t}\;dt\]By parts? ok this is how we would approach it.